# Coupling Electromagnetism to Torsion: Black Holes and Spin-Charge Interactions

### v1

@@ Abstract
The coupling between matter fields and gravity (encoded in the geometry of spacetime) can be realized in various ways. Most commonly, a minimal coupling principle is employed, meaning that all matter fields, except spinors, couple only to the spacetime metric, while spinors additionally couple to the spacetime connection. Non-minimal couplings between matter fields and spacetime curvature can arise, for example, from quantum field theory on curved spacetime through renormalization corrections, in gauge theories of gravity, and in effective field theories. In this article, we consider a non-minimal coupling $F^{\mu\nu}\tilde{R}{\mu\nu}$ between the field strength tensor of the electromagnetic field $F{\mu\nu}$ and the antisymmetric part of the Ricci tensor $\tilde{R}_{[\mu\nu]}$ in Riemann–Cartan geometry, which is based on a general metric-compatible connection with torsion. We find an exact 4-dimensional vacuum solution that generalizes the Reissner–Nordström black hole from Einstein-Maxwell and reveals new interactions between the intrinsic torsion-spin charge and the electric charge. Qualitatively, this solution exhibits two distinct features: the effective charge is not constrained to be positive, and the sign of the electric charge influences its gravitational effects. We also derive slowly rotating solutions in 3 dimensions, representing a generalized slowly rotating BTZ black hole solution with couplings among the magnetic and electric charges, the angular momentum, and the intrinsic torsion-spin charge.
@@ I Introduction
Clearly, pseudo-Riemannian geometry is not the only possibility to describe the geometry of spacetime. On the one hand one can consider the affine connection structure independently from the metric, including properties like torsion Hehl:1976kj ; Bahamonde:2021gfp and non-metricity Hehl:1994ue ; JimenezCano:2021rlu . On the other hand, one could consider non-linear connections as in Finsler Heefer:2024kfi ; Pfeifer:2019wus or Hamilton geometry Barcaroli:2015xda . Since GR and pseudo-Riemmanian geometry together with the standard model or particle physics might not be our final understanding of the gravity, as for example, can be seen from the observational tensions in cosmology CosmoVerse:2025txj and the missing theory of quantum gravity Addazi:2021xuf , spacetime geometries beyond pseudo-Riemannian geometries are investigated as viable extensions of GR. An important aspect in this research is to understand the coupling between matter fields and gravity, i.e. between matter fields and the geometry of spacetime.
@@
In GR, together with the standard model of particle physics, the metric minimal coupling principle is employed, meaning that all matter fields only couple to the spacetime metric, and spinors couple in addition to the Levi-Civita connection. In the framework of quantum field theory on curved spacetime, couplings between curvature and matter fields appear from renormalization corrections to the original action, for example when one considers quantum electrodynamics on curved spacetime DrummondHathrell . It is expected that when one extends the geometry of spacetime beyond pseudo-Riemannian geometry, non-minimal couplings between matter fields and additional geometric fields emerge.
@@
A particularly interesting class of extended spacetime geometries that allow for a broader interaction between spacetime geometry and matter fields, are Riemann-Cartan spacetime geometries. They are based on a spacetime metric and an independent, metric compatible connection with torsion and are the mathematical basis for the formulation of Poincaré gauge gravity theories Hehl:1976kj ; Hehl:1994ue ; Blagojevic:2013xpa ; Obukhov:2022khx . Compared to pseudo-Riemannian spacetime geometry and GR, additional geometric degrees of freedom, describing gravity, are encoded in the torsion tensor. Due to its symmetries, it naturally couples to the electromagnetic field strength tensor, which is why a non-minimal coupling between torsion and electromagnetism has gained particular attention. For example, it has been studied in depth how torsion influences the propagation and structure of electric and magnetic fields Hehl:1999bt ; Itin:2003hr ; deAndrade:1997cj ; Balakin:2007am ; Solanki:2004az ; Hammond:1991ty ; Delhom:2020hkb ; Dias2005 ; Hehl:2016wwp . These studies have demonstrated that torsion can induce anisotropic modifications in the propagation of electromagnetic waves, alter the effective field strength perceived by charged particles, and break the duality symmetry between electric and magnetic components, leading to direction-dependent corrections in electromagnetic interactions. Several works have proposed non-minimal couplings between torsion and the electromagnetic field within effective field theories shapiro2002 ; HehlObukhov2003 , while others have interpreted certain components of the torsion tensor as potential geometric origins of electromagnetic phenomena Hammond2002 . Furthermore, black hole solutions incorporating torsion have been developed in both metric-affine and gauge-theoretic formulations of gravity Lenzen:1985ess ; Tresguerres:1995js ; Tresguerres:1995un ; Vlachynsky:1996zh ; Obukhov:1996pf ; Ho:1997xbz ; Obukhov:1997uc ; Blagojevic2002 ; Baekler:2006de ; Obukhov:2012je ; Giribet:2014hpa ; Cembranos:2016xqx ; Cembranos:2016gdt ; Cembranos:2017pcs ; Obukhov:2019fti ; Bahamonde:2020fnq ; Bahamonde:2020vpb ; Nikiforova:2020sac ; Pfeifer:2021njm ; DAmbrosio:2021zpm ; Bahamonde:2020snl ; Bahamonde:2021qjk ; Bahamonde:2022jue ; Bahamonde:2022esv ; Bahamonde:2021srr ; Bahamonde:2022lvh ; Bahamonde:2022chq ; Bahamonde:2022kwg ; Bahamonde:2022meb ; Gonzalez:2024ifp ; Bahamonde:2024sqo , highlighting the rich phenomenology that torsion introduces to classical gravitational configurations. Also, in lower-dimensional settings, three-dimensional gravity models with electric fields have been constructed and analysed, for example in Blagojevic:2013bu ; Mielke:1991nn ; Mielke:2003xx ; Garcia:2003nm ; Blagojevic:2003wn .
@@
Having identified the action of interest, we derive the corresponding field equations in the tetrad/spin connection formulation of Riemann-Cartan theories of gravity, and explore two physically distinct settings. The first scenario involves a static, spherically symmetric configuration in four dimensions, where we find exact black hole solutions in which torsion plays a central role. Not only does it alter the geometry of spacetime, but it also modifies the effective electric charge through the emergence of an intrinsic spin charge that couples directly to the electric field—giving rise to corrections absent in the standard Reissner–Nordström framework. For the second scenario, we turn to three-dimensional gravity with a negative cosmological constant, a setting that has been extensively explored in the context of generalizations of the BTZ black hole Banados:1992wn ; Banados:1992gq , with torsion acting as a cosmological constant Mielke:2003xx ; Garcia:2003nm ; Blagojevic:2003wn ; Canfora:2007fw ; Blagojevic:2008ip ; Blagojevic:2008xz ; Cvetkovic:2008zz ; Vasquez:2009mk ; Blagojevic:2009xw ; Cvetkovic:2018ati ; Canfora:2010rh ; Blagojevic:2013aaa ; Becar:2013qba ; Gonzalez:2014pwa ; Ma:2014tka ; Aviles:2023igk . For the theory under discussion we construct exact solutions in the regime of slow rotation and the torsion tensor acts as an electric potential. In the static limit, the coupling between spin and electric charge reflects the behaviour observed in four dimensions. Remarkably, once rotation is introduced perturbatively, a new term emerges that links the intrinsic spin charge to the angular momentum, an effect absent in the original charged BTZ geometry Martinez:1999qi . Altogether, our findings reveal new mechanisms through which torsion can influence gravitational and gauge dynamics, leading to nontrivial spacetime structures and couplings with physical observables such as electric charge and angular momentum.
We organise the presentation of our results as follows. In Sec.II, we present the basic concepts of Riemann-Cartan geometries. Then, in Sec.III, we construct and introduce our theory with new couplings between the electromagnetic field and the generalised Ricci tensor. In Secs. IV and V, we present new black hole solutions in four and three dimensions, respectively. Finally, in Sec. VI, we summarise the main conclusions and outline potential directions for future research within this framework. Throughout this article we use the following conventions: The signature of the metric is mostly minus (+,,,)(+,-,...,-)( + , - , … , - ). A tilde $\tilde{}$ over an object denotes quantities computed with the general metric compatible connection with torsion.
@@ II Riemann-Cartan geometries
A Riemann-Cartan spacetime $(M,g,\tilde{\Gamma})$ is an n-dimensional manifold $M$ equipped with a Lorentzian metric and a metric compatible connection $\tilde{\Gamma}$ with torsion Hehl:1994ue ; Obukhov:2022khx .
@@
which allows us to decompose the $\tilde{\Gamma}$ covariant derivative ~$\tilde{\nabla}_{\mu}v^{\lambda}$ of an arbitrary vector $v^{\lambda}$ into a Levi-Civita covariant derivative $\nabla_{\mu}v^{\lambda}$ and the so-called contortion tensor $K^{\lambda}\,_{\rho\mu}$
~$\tilde{\nabla}_{\mu}v^{\lambda}$ $=\nabla_{\mu}v^{\lambda}+K^{\lambda}\,_{\rho\mu}v^{\rho}$ (2)
@@
[~$[\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\,v^{\lambda}=\tilde{R}^{\lambda}\,_{\rho\mu\nu}\,v^{\rho}+T^{\rho}\,_{\mu\nu}\,\tilde{\nabla}_{\rho}v^{\lambda}\,,$ (5)
@@
For constructing Riemann-Cartan theories of gravity and to introduce spinors, it is convenient to introduce tetrads $e^{a}{}_{\mu}$ and a principal bundle connection $\tilde{\omega}_{\mu}\in\mathfrak{so}(1,3)$ that leads to a Poincaré gauge characterisation of the geometry of spacetime Hehl:1976kj ; ponomarev2017gauge . The metric and the affine connection can be related to the tetrad and spin connection as
@@ III Coupling electromagnetic fields with torsion on Riemann-Cartan spacetime
where the electromagnetic field strength is given in terms of the electromagnetic potential $1$-form $A$ as
@@
The $\tilde{R}^{2}$ terms guarantees that the degrees of freedom of $\tilde{\Gamma}$ that generate torsion are also propagating, as, when expanded with help of equation (10), it generates terms of the form $\nabla T\nabla T$. In particular the spin-0 part of torsion propagates consistent with the causal structure of spacetime Yo:1999ex .
@@
In the literature, also, additional quadratic terms in the torsion tensor (mass terms for the torsion), are often considered Bahamonde:2021gfp ,
@@
While the direct torsion-field strength ($TTF$) couplings $k_{4},k_{5}$ and $k_{6}$ have been introduced and studied for example in Itin:2003hr , a detailed study of the $\tilde{R}F$ field strength Riemann-Cartan curvature coupling is missing. We will provide this study in the following.
@@ IV Exact spherically symmetric black hole solutions in 4 dimensions
$T^{\vartheta}\,_{t\varphi}$ $=T^{\varphi}\,_{\vartheta t}\sin^{2}\vartheta=t_{5}(r)\sin{\vartheta}\,,\quad T^{\vartheta}\,_{r\varphi}=T^{\varphi}\,_{\vartheta r}\sin^{2}\vartheta=t_{6}(r)\sin{\vartheta}\,,$ (36)
@@ IV.2 Solution with an additional kinetic term for torsion
$t_{2}(r)$ =±$=\pm\sqrt{\Psi_{1}(r)\Psi_{2}(r)}\left(t_{4}(r)-t_{1}(r)-\frac{A_{0}^{\prime\prime}(r)}{A_{0}^{\prime}(r)}\right)\pm\frac{\Psi_{2}(r)\Psi_{1}^{\prime}(r)-\Psi_{1}(r)\Psi_{2}^{\prime}(r)}{2\sqrt{\Psi_{1}(r)\Psi_{2}(r)}}\,,$ (41)
@@
where $c_{1}$ and $\kappa_{s}$ are constants that absorb combinations of the parameters in the theory. In particular, the intrinsic spin charge $\kappa_{s}$ is identified through the term $1/(r\Psi(r))$, since the axial component of the torsion, $S_{\mu}$, behaves like a Coulomb potential that determines the torsion tensor (see also Bahamonde:2020fnq ), and is given by,
@@
As we assumed $k_{2}\neq 0$ in our computations, the standard Einstein-Maxwell case would be recovered by imposing $k_{3}=0$. This metric is of Reissner–Nordström form but exhibits a few important differences. The interaction between the spin charge $\kappa_{\rm s}$ and the electric charge $q$ can, depending on their magnitudes, change the sign of the term in brackets, which would completely alter the resulting phenomenology of the spacetime geometry. Further, the effective charge is not restricted to always be positive as in the standard Einstein-Maxwell case. Then, in principle, one can avoid the Cauchy horizon Poisson:1989zz ; Hollands:2019whz ; Maeda:2005yd with these corrections, and it has also been pointed out to occur in non-linear electrodynamics Hale:2025ezt .
@@
±$\pm\frac{1}{\Psi(r)}\Big{[}\frac{r^{2}\left(r\Psi^{\prime}(r)+\Psi(r)-1\right)}{k_{3}q}+\left(\frac{k_{1}}{k_{3}}+\frac{k_{3}}{32k_{2}}\right)q+2rt_{4}(r)\kappa_{s}+2\kappa_{s}\Big{]}\,,$

### v2

@@ Abstract
The coupling between matter fields and gravity (encoded in the geometry of spacetime) can be realized in various ways. Most commonly, a minimal coupling principle is employed, meaning that all matter fields, except spinors, couple only to the spacetime metric, while spinors additionally couple to the spacetime connection. Non-minimal couplings between matter fields and spacetime curvature can arise, for example, from quantum field theory on curved spacetime through renormalization corrections, in gauge theories of gravity, and in effective field theories. In this article, we consider a non-minimal coupling $F^{\mu\nu}\tilde{R}_{\mu\nu}$ between the field strength tensor of the electromagnetic field $F_{\mu\nu}$ and the antisymmetric part of the Ricci tensor $\tilde{R}_{[\mu\nu]}$ in Riemann–Cartan geometry, which is based on a general metric-compatible connection with torsion. We find an exact 4-dimensional vacuum solution that generalizes the Reissner–Nordström black hole from Einstein-Maxwell and reveals new interactions between the intrinsic torsion-spin charge and the electric charge. Qualitatively, this solution exhibits two distinct features: the effective charge is not constrained to be positive, and the sign of the electric charge influences its gravitational effects. We also derive slowly rotating solutions in 3 dimensions, representing a generalized slowly rotating BTZ black hole solution with couplings among the magnetic and electric charges, the angular momentum, and the intrinsic torsion-spin charge.
@@ I Introduction
Clearly, pseudo-Riemannian geometry is not the only possibility to describe the geometry of spacetime. On the one hand one can consider the affine connection structure independently from the metric, including properties like torsion Hehl et al. (1976); Bahamonde et al. (2023b) and non-metricity Hehl et al. (1995); Jiménez Cano (2021). On the other hand, one could consider non-linear connections as in Finsler Heefer (2024); Pfeifer (2019) or Hamilton geometry Barcaroli et al. (2015). Since GR and pseudo-Riemmanian geometry together with the standard model or particle physics might not be our final understanding of the gravity, as for example, can be seen from the observational tensions in cosmology Di Valentino and others (2025) and the missing theory of quantum gravity Addazi and others (2022), spacetime geometries beyond pseudo-Riemannian geometries are investigated as viable extensions of GR. An important aspect in this research is to understand the coupling between matter fields and gravity, i.e. between matter fields and the geometry of spacetime.
@@
In GR, together with the standard model of particle physics, the metric minimal coupling principle is employed, meaning that all matter fields only couple to the spacetime metric, and spinors couple in addition to the Levi-Civita connection. In the framework of quantum field theory on curved spacetime, couplings between curvature and matter fields appear from renormalization corrections to the original action, for example when one considers quantum electrodynamics on curved spacetime Drummond and Hathrell (1980). It is expected that when one extends the geometry of spacetime beyond pseudo-Riemannian geometry, non-minimal couplings between matter fields and additional geometric fields emerge.
@@
A particularly interesting class of extended spacetime geometries that allow for a broader interaction between spacetime geometry and matter fields, are Riemann-Cartan spacetime geometries. They are based on a spacetime metric and an independent, metric compatible connection with torsion and are the mathematical basis for the formulation of Poincaré gauge gravity theories Hehl et al. (1976, 1995); Blagojević and Hehl (2013); Obukhov (2023). Compared to pseudo-Riemannian spacetime geometry and GR, additional geometric degrees of freedom, describing gravity, are encoded in the torsion tensor. Due to its symmetries, it naturally couples to the electromagnetic field strength tensor, which is why a non-minimal coupling between torsion and electromagnetism has gained particular attention. For example, it has been studied in depth how torsion influences the propagation and structure of electric and magnetic fields Hehl and Obukhov (2001); Itin and Hehl (2003); de Andrade and Pereira (1999); Balakin et al. (2008); Solanki et al. (2004); Hammond (1991); Delhom (2020); Dias and Moraes (2005); Hehl et al. (2016). These studies have demonstrated that torsion can induce anisotropic modifications in the propagation of electromagnetic waves, alter the effective field strength perceived by charged particles, and break the duality symmetry between electric and magnetic components, leading to direction-dependent corrections in electromagnetic interactions. Several works have proposed non-minimal couplings between torsion and the electromagnetic field within effective field theories Shapiro (2002); Hehl and Obukhov (2003), while others have interpreted certain components of the torsion tensor as potential geometric origins of electromagnetic phenomena Hammond (2002). Furthermore, black hole solutions incorporating torsion have been developed in both metric-affine and gauge-theoretic formulations of gravity Lenzen (1985); Tresguerres (1995b, a); Vlachynsky et al. (1996); Obukhov et al. (1996); Ho et al. (1997); Obukhov and Hehl (1998); Blagojević (2002); Baekler and Hehl (2006); Obukhov and Hehl (2012); Giribet et al. (2014); Cembranos et al. (2017); Cembranos and Valcarcel (2017); Cembranos and Gigante Valcarcel (2018); Obukhov (2019); Bahamonde and Valcarcel (2020); Bahamonde and Pfeifer (2021); Nikiforova and Damour (2020); Pfeifer and Schuster (2021); D’Ambrosio et al. (2022); Bahamonde et al. (2021); Bahamonde and Valcarcel (2022); Bahamonde et al. (2022b, c, d, a, 2023c, 2023a, e); González et al. (2025); Bahamonde and Gigante Valcarcel (2024), highlighting the rich phenomenology that torsion introduces to classical gravitational configurations. Also, in lower-dimensional settings, three-dimensional gravity models with electric fields have been constructed and analysed, for example in Blagojevic et al. (2013); Mielke and Baekler (1991); Mielke and Rincon Maggiolo (2003); Garcia et al. (2003); Blagojevic and Vasilic (2003).
@@
Having identified the action of interest, we derive the corresponding field equations in the tetrad/spin connection formulation of Riemann-Cartan theories of gravity, and explore two physically distinct settings. The first scenario involves a static, spherically symmetric configuration in four dimensions, where we find exact black hole solutions in which torsion plays a central role. Not only does it alter the geometry of spacetime, but it also modifies the effective electric charge through the emergence of an intrinsic spin charge that couples directly to the electric field—giving rise to corrections absent in the standard Reissner–Nordström framework. For the second scenario, we turn to three-dimensional gravity with a negative cosmological constant, a setting that has been extensively explored in the context of generalizations of the BTZ black hole Banados et al. (1992, 1993), with torsion acting as a cosmological constant Mielke and Rincon Maggiolo (2003); Garcia et al. (2003); Blagojevic and Vasilic (2003); Canfora (2009); Blagojevic and Cvetkovic (2008b, a); Cvetkovic and Blagojevic (2008); Vasquez (2011); Blagojevic et al. (2009); Cvetković and Simić (2019); Canfora and Giacomini (2010); Blagojević et al. (2013); Becar et al. (2014); González et al. (2015); Ma et al. (2014); Avilés et al. (2023). For the theory under discussion we construct exact solutions in the regime of slow rotation and the torsion tensor acts as an electric potential. In the static limit, the coupling between spin and electric charge reflects the behaviour observed in four dimensions. Remarkably, once rotation is introduced perturbatively, a new term emerges that links the intrinsic spin charge to the angular momentum, an effect absent in the original charged BTZ geometry Martinez et al. (2000). Altogether, our findings reveal new mechanisms through which torsion can influence gravitational and gauge dynamics, leading to nontrivial spacetime structures and couplings with physical observables such as electric charge and angular momentum.
We organise the presentation of our results as follows. In Sec.II, we present the basic concepts of Riemann-Cartan geometries. Then, in Sec.III, we construct and introduce our theory with new couplings between the electromagnetic field and the generalised Ricci tensor. In Secs. IV and V, we present new black hole solutions in four and three dimensions, respectively. Finally, in Sec. VI, we summarise the main conclusions and outline potential directions for future research within this framework. Throughout this article we use the following conventions: The signature of the metric is mostly minus (+,,,)(+,-,...,-). A tilde $\tilde{}$ over an object denotes quantities computed with the general metric compatible connection with torsion.
@@ II Riemann-Cartan geometries
[…]
@@
which allows us to decompose the $\tilde{\Gamma}$ covariant derivative $\tilde{\nabla}_{\mu}v^{\lambda}$ of an arbitrary vector $v^{\lambda}$ into a Levi-Civita covariant derivative $\nabla_{\mu}v^{\lambda}$ and the so-called contortion tensor $K^{\lambda}\,_{\rho\mu}$
$\tilde{\nabla}_{\mu}v^{\lambda}$ $=\nabla_{\mu}v^{\lambda}+K^{\lambda}\,_{\rho\mu}v^{\rho}$ (2)
@@
$\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\,v^{\lambda}=\tilde{R}^{\lambda}\,_{\rho\mu\nu}\,v^{\rho}+T^{\rho}\,_{\mu\nu}\,\tilde{\nabla}_{\rho}v^{\lambda}\,$, (5)
@@
For constructing Riemann-Cartan theories of gravity and to introduce spinors, it is convenient to introduce tetrads $e^{a}{}_{\mu}$ and a principal bundle connection $\tilde{\omega}_{\mu}\in\mathfrak{so}(1,3)$ that leads to a Poincaré gauge characterisation of the geometry of spacetime Hehl et al. (1976); V. N. Ponomarev, A. O. Barvinsky, and Yu. N. Obukhov (2017). The metric and the affine connection can be related to the tetrad and spin connection as
@@ III Coupling electromagnetic fields with torsion on Riemann-Cartan spacetime
where the electromagnetic field strength is given in terms of the electromagnetic potential $1$-form AA as
@@
The term controlled by $k_{2}$ may effectively behave as a kinetic contribution for torsion (see (10)) when suitable torsional invariants are incorporated.
@@
In the literature, also, additional quadratic terms in the torsion tensor (mass terms for the torsion), are often considered Bahamonde et al. (2023b),
@@
While the direct torsion-field strength ($TTF$) couplings $k_{4},k_{5}$ and $k_{6}$ have been introduced and studied for example in Itin and Hehl (2003), a detailed study of the $\tilde{R}F$ field strength Riemann-Cartan curvature coupling is missing. We will provide this study in the following.
@@
It is worth emphasising that, in a strict effective-field-theory (EFT) sense, the torsion–electromagnetism coupling introduced in Eq. (26) is not mandated solely by diffeomorphism and $U(1)$ invariance. Our goal here, however, is not to construct a complete EFT expansion, but rather to explore a specific geometric realisation of non-minimal couplings within the Riemann–Cartan framework. This point of view follows a long-standing tradition in torsion electrodynamics (see, e.g., Itin and Hehl (2003); Hehl and Obukhov (2001); de Andrade and Pereira (1999); Balakin et al. (2008); Solanki et al. (2004); Hammond (1991); Delhom (2020); Dias and Moraes (2005); Hehl et al. (2016); Shapiro (2002); Hehl and Obukhov (2003)), where the emphasis is placed on covariant and geometrically motivated operators, rather than on EFT completeness. Accordingly, the operators in Eq. (26) should be regarded as a representative, geometrically natural subset of all possible non-minimal interactions, capturing the essential phenomenology of the torsionful electromagnetic sector.
@@ IV Exact spherically symmetric black hole solutions in 4 dimensions
[…]
@@ IV.2 Solution with an additional kinetic term for torsion
$t_{2}(r)$ $=\pm\sqrt{\Psi_{1}(r)\Psi_{2}(r)}\left(t_{4}(r)-t_{1}(r)-\frac{A_{0}^{\prime\prime}(r)}{A_{0}^{\prime}(r)}\right)\pm\frac{\Psi_{2}(r)\Psi_{1}^{\prime}(r)-\Psi_{1}(r)\Psi_{2}^{\prime}(r)}{2\sqrt{\Psi_{1}(r)\Psi_{2}(r)}}\,$, (41)
@@
where $c_{1}$ and $\kappa_{s}$ are constants that absorb combinations of the parameters in the theory. In particular, the intrinsic spin charge $\kappa_{s}$ is identified through the term $1/(r\Psi(r))$, since the axial component of the torsion, $S_{\mu}$, behaves like a Coulomb potential that determines the torsion tensor (see also Bahamonde and Valcarcel (2020)), and is given by,
@@
As we assumed $k_{2}\neq 0$ in our computations, the standard Einstein-Maxwell case would be recovered by imposing $k_{3}=0$. This metric is of Reissner–Nordström form but exhibits a few important differences. The interaction between the spin charge $\kappa_{\rm s}$ and the electric charge $q$ can, depending on their magnitudes, change the sign of the term in brackets, which would completely alter the resulting phenomenology of the spacetime geometry. Further, the effective charge is not restricted to always be positive as in the standard Einstein-Maxwell case. Then, in principle, one can avoid the Cauchy horizon Poisson and Israel (1989); Hollands et al. (2020); Maeda et al. (2005) with these corrections, and it has also been pointed out to occur in non-linear electrodynamics Hale et al. (2025).
@@
$\pm\frac{1}{\Psi(r)}\Big[\frac{r^{2}\left(r\Psi^{\prime}(r)+\Psi(r)-1\right)}{k_{3}q}+\left(\frac{k_{1}}{k_{3}}+\frac{k_{3}}{32k_{2}}\right)q+2rt_{4}(r)\kappa_{s}+2\kappa_{s}\Big]\,$,